OpenStudy (anonymous):

Okay... Im having the hardest time with understanding these... I just cant do it! Write the equation of the line with a slope of -1/3 and passing through the point (6, -4). Write the equation of the line passing through the points (0, -4) and (-2, 2). Write the equation of the line passing through the points (-6,1) and (-4,2). Write the equation of the line with an undefined slope, passing through the point (2, 5). Hint: refer back to lesson 4.04 for help with this one. Write the equation of the line on the following graph.

OpenStudy (anonymous):

An equation of a line will have the form y = Ax + B so your job is to find A, B. Slope of -1/3 means A = -1/3. Then you plug the pair into y = -1/3 x + B to find B Line passing thru two points: you plug the two pairs into the form y = Ax + B to get two equations and then solve the system of equations. So many lines can be drawn when only 1 point is given, so we'll need to look into that hint.